How much energy does the refrigerator use in a day, in kWh? What is the total amount
of energy saved in a year by choosing the more efficient appliance?
Answer #1

The power consumption of domestici referigerators typically between 100 and 200 watts.

Over a full day they are likely to use around 1 to 2 kilowatt-hours (kWh)

Lets suppose power ratings of refrigerator is 180 watt(W) and in a day it is used for 8 hours

Therefore energy consumed in a day (E) = power * time = 180 * 8= 1440 Wh

In kWh ,. E = 1440 /1000 = 1.440 kWh  

So , energy consumption of this conventional refrigerator in a year is E1 = E * 365 = 525.6 kWh

In this modern era the power ratings of efficient refrigerator is about 100 W

So , energy consumption of efficient refrigerator is E2 = 0.100*8*365=292 kWh

Energy can be saved in a year = E1-E2 =525.6-292=233.6 kWh